OpenStudy (anonymous):

Find the point of intersection between: Pt=(3t-4,2t-1) and 4x+3y=18. Best answer gets medal!

ganeshie8 (ganeshie8):

what do you know about parametric form of line ?

OpenStudy (anonymous):

In y=mx+b form it is y=2/3x + 5/3

OpenStudy (triciaal):

one approach start with a sketch of the line when x = 0 y = 6 when y = 0 x =18/4 = 9/2

ganeshie8 (ganeshie8):

Excellent! so you have eliminated the parameter already you have two equations : \[y = \frac{2}{3}x + \frac{5}{3}\tag {1}\] \[4x+3y = 18 \tag{2}\] solve them using substitution

OpenStudy (anonymous):

and 4x + 3y = 18 in y=mx+b is y=-4/3 + 6

OpenStudy (anonymous):